A Ferris wheel with radius 14.0 m is turning about a horizontal axis through its center. The linear speed of a passenger on the rim is constant and equal to 7.00 m/s. What are the magnitude and direction of the passenger’s acceleration as she passes through (a) the lowest point in her circular motion? (b) The highest point in her circular motion? (c) How much time does it take the Ferris wheel to make one revolution?


1
Expert's answer
2021-12-15T16:22:17-0500

a)

"a_1=g+\frac{v^2}r=13.3~\frac m{s^2}," down,

b)

"a_1=g-\frac{v^2}r=6.3~\frac{m}{s^2}," down,

c)

"T=\frac{2\pi}{\omega}=\frac{2\pi r}v=12.6~s."